In a random sample of 19 people, the mean commute time to work was 31.7 minutes and the standard deviation was 7.2 minutes. Assume the population is normally distributed and use a​ t-distribution to construct a 80​% confidence interval for the population mean

What is the margin of error of μ​?

Interpret the results.

The confidence interval for the population mean μ is (,)
Answer #1

Solution :

Given that,

sample size = n = 19

Degrees of freedom = df = n - 1 = 19 - 1 = 18

t /2,df = 1.330

Margin of error = E = t/2,df * (s /n)

= 1.330 * (7.2 / 19)

Margin of error = E = 2.2

The 80% confidence interval estimate of the population mean is,

- E < < + E

31.7 - 2.2 < < 31.7 + 2.2

29.5 < < 33.9

The confidence interval for the population mean μ is (29.5, 33.9)